Please help me answer this problem:
Worksheet says: Solve each system by substitution. check your solution.
a-1.2b = -3
0.2b+ 0.6a= 12

Answer by Osran_Shri(18)

here we have to find out values of "a" and "b" by substitution method and verify the solution.
a-1.2b = -3 ...............Equation-1
0.2b+ 0.6a= 12 ...............Equation-2
From Equation-1:
a = (-3 + 1.2b )= (1.2b - 3)...Equation-3
Now substitute "a" from above (Equation-3) into Equation-2 to get following equation:
0.2b + 0.6(1.2b-3)=12
0.2b + 0.72b - 1.8 = 12
0.92b - 1.8 = 12
0.92b = 12 + 1.8 = 13.8
b = 13.8/0.92 = 15
Therefore b = 15 Substitute b=15 into Equation-3 to get "a" as below:
a = 1.2b - 3 = 1.2*15 - 3 = 18 - 3 = 15
Hence values of : a=15 & b=15
Verification of Solution:
--------------------------
Substitute values of a & b into Equation-1 to get -3 as below:
a-1.2b = 15 - 1.2*15 = 15 - 18 = -3 --------> a-1.2b= -3
Now substitute values of a & b into Equation-2 to get 12 as below:
0.2b+ 0.6a = 0.2*15 + 0.6*15 = 3 + 9 = 12 -------> 0.2b+0.6a=12
thats it....
B.Shridhar